5) you are required to make agar plates for your microbiology lab. You must prepare 100 plates with 20 ml of media per plate. to solidify the media, you will add 1.5% agar. how will you prepare the lab?
Answer #1

we have to prepare 100 plate of media,each plate require 20 ml of media and we use 1.5% agar for solidify media

so that

total media we have to prepare = 100 x 20 = 2000ml

total agar required = 2000 x 1.5 / 100 = 3000/100= 30gram

to prepare lab we have to prepare 2000 ml of media and add 30 gram of agar powder in media, after that autoclave the media and clean plates, pour media in plate aseptically and keep plate to solidify in aseptic condition.